OpenStudy (anonymous):

The area of a triangle is 45.5 square meters and the base is 10 meters. What is the height of this triangle? A. 9.1 m B. 455 m C. 4.55 m D. 227.5 m

OpenStudy (anonymous):

The formula for the area of a triangle is A=1/2bh. This means that one half of the base times height is the area of the triangle. since you don't the height, fill in the info that you do know: 45.5=1/2(10)h or 45.5=5h Now, you should have no problem solving for h :)

OpenStudy (anonymous):

answer is 9.1 m